OpenStudy (anonymous):

a vine 6 cm tall is planted in a garden. it grows 4 centimeters a day. how many days will it take to reach 122 cm in height? a. 32 b. 30.5 c. 29 d. 27

OpenStudy (radar):

Let x = the number of days. then \ 6+4x=122 cm 4x=122-6=116 x=29

OpenStudy (anonymous):

thank you , you're really helping me! :)

OpenStudy (radar):

U R Welcome